how to
factor q2-3q+18=0

Answer by RAY100(1637)

q^2 -3q +18 =0
.
if it was q^2 -3q-18=(q-6)(q+3)
.
however as it stands use quadratic formula
.
a=1,,,b=-3,,,c=18
.
q= [ -(-3) +/- sqrt{ (-3)^2 -4(1)(18) ] / 2(1)
.
q= [3 +/- sqrt{9-72} ] /2
.
q=[ 3+/- (7.94 i ) /2
.
q= 3/2 +/- 3.97 i